OpenStudy (anonymous):

Instead of doing implicit differentiation, can you solve for y and then take the derivative?

OpenStudy (anonymous):

for example, 2 - 8xy = 5x - 3y 3y -8xy = 5x - 2 y = (5x-2) / (3-8x), dy/dx = 5(3-8x) - (-8)(5x-2) / (3-8x)^2

hero (hero):

You can, but only when you are able to solve for y. You cannot with functions where you are not able to solve for y

OpenStudy (anonymous):

This is right I think...

OpenStudy (anonymous):

what is an example of a function you cannot solve for y?

hero (hero):

\[x^2 - 2y^3 +4y = 2\]

OpenStudy (anonymous):

so functions like this must have y on the other side of dy/dx

OpenStudy (anonymous):

so you need implicit differentiation because sometimes you cannot single out the y?

hero (hero):

Precisely

OpenStudy (anonymous):

and most of the time, you cannot solve for y because they are not functions...